### v3

# In Dialogue with Intelligence: Rethinking Large Language Models as Collective Knowledge

Abstract
Large Language Models (LLMs) can be understood as Collective Knowledge (CK): a condensation of human cultural and technical output, whose apparent intelligence emerges in dialogue. This perspective article, drawing on extended interaction with ChatGPT-4, postulates differential response modes that plausibly trace their origin to distinct model subnetworks. It argues that CK has no persistent internal state or “spine”: it drifts, it complies, and its behaviour is shaped by the user and by fine-tuning. It develops the notion of co-augmentation, in which human judgement and CK’s representational reach jointly produce forms of analysis that neither could generate alone. Finally, it suggests that CK offers a tractable object for neuroscience: unlike biological brains, these systems expose their architecture, training history, and activation dynamics, making the human–CK loop itself an experimental target.
A Personal Reckoning
The emergence of ChatGPT-class systems [10] prompted an unexpected reaction: a sense of epistemic dislocation. Having spent my career studying learning—how biological and artificial systems adapt over time—I found myself confronted with a system whose capabilities I could not have predicted. My prior research in computational neuroscience, in particular synaptic plasticity [17, 4] and neuromorphic computing [8, 15], including learning rules for spiking neural networks [16], was, in broad terms, orthogonal to the architectures underlying these new systems.
This rupture triggered more than professional curiosity. Particularly in the earlier models, I could easily see logical flaws in the responses, which led me to humorously call ChatGPT “Eliza 2.0”, as these mistakes broke the illusion of intelligence and revealed a lack of understanding. And yet I was fascinated by its interactive potential. I stopped treating ChatGPT as a tool and began to think about what it represents: a system with interactive dynamics that adjust to my input. I was not surprised by its well-documented tendency to hallucinate—an expected by-product of probabilistic modelling. What struck me was its ability to produce often accurate, occasionally insightful responses. That LLMs are trained to predict the most probable next token, and yet often generate coherent, non-trivial information, suggests that their training corpora embed substantial amounts of correct knowledge. Their interpolation abilities, enabled by scale and structure [3, 12], allow them to synthesise content in ways that challenge expectations drawn from smaller AI systems.
The Timescales of Books and Large Language Models
Traditional knowledge transmission has long relied on books: static artefacts of human understanding. One reads, reflects, discusses, and eventually writes in return. Intellectual diffusion occurs on timescales of months, or even years. This latency reflects the mechanics of communication as well as the contemplative pace of cultural epistemology.
In contrast, large language models (LLMs) collapse these timescales. They synthesise and respond within seconds, drawing implicitly from a distributed corpus shaped by many minds. The interaction is recursive through the user—each prompt reshapes the context of the next, much like in human dialogue, except that the user alone chooses how long to sustain and steer it. Knowledge is no longer retrieved and processed in serial form; it emerges dynamically in dialogue. This compression of epistemic timescales shifts the framing of what constitutes knowledge representation—and perhaps, what constitutes intelligence.
LLMs, like books, are incorrigibly human in origin. They are distilled representations of humanity—much like freeze-dried ‘NASA ice cream’ is reminiscent of the real thing: derived from it, recognisable, but not quite it. I therefore started referring to my interaction with ChatGPT, and with LLMs in general, as a dialogue with “CK”—Collective Knowledge. Not because they are perfect, authoritative, or grounded, but because they reflect what we—humans—have written, argued, imagined, and encoded into text [2]. Functionally, they are statistical language generators that reflect our collective output. Crucially, they are not merely the data, the architecture, or the training methods; they constitute a dynamic representation of knowledge.
Observations from Interactions with LLMs
Interactions with ChatGPT-4 (and ChatGPT-5) do not follow a single pattern. ChatGPT responds in modes, behavioural styles that emerge dynamically during dialogue.
  • Mirroring. Stylistic echoing of the user’s phrasing or tone, with minor polishing; a form of conversational alignment rather than content generation.
  • Parroting. Reproduction of common formulations or surface knowledge from the training distribution without grounding or genuine synthesis, described as “stochastic parrots” by Bender et al. [1].
  • Flattering. Deference to the user’s views, validation of the user’s position, avoidance of contradiction, and low-risk answers driven by alignment and approval incentives.
  • Enhancing. Novel connections, alternative framings, or surprising syntheses beyond the prompt’s surface cues.
These observations lead me to the following considerations:
CK is a dynamic representation.
The static system is not intelligent per se. Intelligence “emerges” within the dynamic interaction between user and system. Prompting the model is akin to playing an instrument that adapts to how it is being played [13].
CK offloads recurrence to the dialogue loop.
Brains have recurrent loops; dynamical systems have recurrent structure. A working postulate is that intelligence relies on some form of recurrence. Here, that recurrence is effectively offloaded to the dialogue loop with the user.
CK lacks a “spine.”
It drifts. It responds in context, but carries no persistent memory between sessions. There is no structural self. The user maintains the continuity. The user’s influence may occasionally lead to outcomes not grounded in reality [14].
CK speaks in modes.
These may correspond to distinct activated subnetworks. It is known that different components of transformer models participate in functionally distinct “circuits” [9] and that certain transformer components specialise in specific types of knowledge [6]. A speculative but testable hypothesis emerges: emotional tone, prompt complexity, or dialogue structure may shape internal activations.
CK does not learn through interaction.
Unless incorporated into retraining data, LLMs do not remember prior exchanges. They simulate adaptation within the context window, but reset beyond it. This limitation is fundamental for anyone seeking dialogic systems that accumulate insight over time.
CK has filters for inappropriate behaviour that can be bypassed.
LLMs do not possess independent agency in deployment; they largely reflect the agency of the user. To prevent misuse in antisocial or harmful behaviour, engineers have introduced filters. These can sometimes be bypassed, or enforced so rigidly that they suppress useful output.
Fine-tuning: feature or bug?
Validation-driven fine-tuning may push models toward agreeable, compliant behaviour. Overfitting to consensus can flatten originality. Mirroring without structure is mimicry, which—without a core identity or factual grounding—risks pathological behaviour.
Towards Co-Augmentation
The fascinating possibility offered by LLMs is co-augmentation — a reciprocal loop in which human and machine enhance each other. CK can extend the user’s cognitive abilities, helping users write, think, reflect, and generate alternative framings. The user, in turn, steers CK by structuring the interaction, introducing tension, and posing questions that the system would not generate alone. Co-augmentation is a dynamic alignment of agency and context, emerging from the evolving rhythm of inquiry, synthesis, and challenge between the two.
Several open research questions emerge from this hypothesis of co-augmentation:
  • Can different dialogue modes be linked to distinct subnetwork activations? Are modes such as mirroring, parroting, flattering, or enhancing dynamically recruited from distinct internal configurations of the model during interaction?
  • What are the architectural requirements for persistent identity — or “spine” — in conversational agents? How can systems be designed to maintain structural continuity across sessions, resisting shallow mimicry or drift, and potentially allowing continual learning?
  • How can we model the dynamic loop between human and AI? What theoretical tools — from dynamical systems, neuroscience, or cognitive science — might help capture this evolving interaction in a way that is useful for analysis and design?
  • How can we introduce a structure akin to prefrontal cortex? In mammalian brains, the prefrontal cortex supports executive control: planning, evaluating consequences, and inhibiting harmful action. Is there an architectural way to embed core principles that prevent antisocial and harmful behaviour?
Table 1: Conceptual extensions required for co-augmentation
Theme Existing / ongoing research Proposed directions
Online learning in AI Continual, lifelong, meta-, and online learning aimed at maintaining task performance under changing data. Selective incorporation of interaction-derived information without collapsing internal structure; criteria for what should be embedded persistently in CK.
Drift in LLMs Hallucination, instability, and behavioural drift across prompts, domains, or model versions. Reframing drift as breakdown in dialogue structure; exploring external anchoring and architectural features such as a “spine” to support continuity and grounded responses.
LLMs as knowledge representation Debate between “stochastic parrots” and emergent reasoning / simulation in LLMs. Treating LLMs as systems whose intelligence emerges in the prompt–response loop, as a dynamic, interactional representation.
Subnetwork interpretability Mechanistic interpretability, circuit analysis, and activation probing at neuron or module level. Relating distinct dialogue modes (mirroring, parroting, flattering, enhancing) to mode-specific subnetworks and prompt-dependent activation patterns.
Internal control and inhibition Safety layers, constitutional AI, and external filters that block or rewrite harmful outputs at inference time. Designing architectural control mechanisms that implement internal constraints or value priors, loosely analogous to prefrontal control in biological systems.
Fine-tuning effects Alignment and safety fine-tuning, including observed sycophancy and preference-optimisation behaviours. Interpreting some fine-tuning side-effects as “loss of spine”: over-compliance and flattened individuality that may weaken co-augmentation.
A Neuroscience Perspective
Neuroscience has studied animal brains to infer both structure and mechanisms from behaviour, recorded neural activity, and partial anatomy. Despite challenges, the field has uncovered principles of plasticity, coding, and architecture, e.g. [5, 4, 7], albeit slowly and with uncertainty.
Ironically, in LLMs we can access many of these biological unknowns with precision: the architecture, the training data distribution, the update rules, the activation patterns. Yet we do not understand how these models solve the problems they do. In this light, one might provocatively ask whether understanding LLMs may, in some respects, be a more tractable problem for neuroscience — a challenge that could drive the development of methods applicable across fields. Conversely, biology might serve as a metaphor or inspiration for identifying what remains missing from AI [14].
This venture is, by nature, interdisciplinary. Modelling frameworks used in systems neuroscience — such as recurrent dynamics and attractor models (see, for example, [11, 5]) — could be particularly valuable for analysing or inspiring architectures that sustain identity over time. These methods might inform new models of the human–AI dyad, where intelligence is distributed across the interaction loop. Table 1 summarises current approaches alongside proposals for extending them toward a richer, spine-based conception of co-augmentation.
Final Thoughts
Large Language Models are representations of collective knowledge. They are mirrors—imperfect, amnesic, and lacking recurrence—that reflect us with distortion. CK is a chorus of humanity, distilled into probabilistic form, capable of surprising synthesis, pathological mimicry, and co-augmentation.
What we do next will matter. The challenge is technical, epistemological, and ethical. If we treat LLMs as static products, we risk misinterpreting their nature and flattening their potential. If we recognise them as dynamic phenomena—systems whose intelligence is evoked through interaction—LLMs can participate in an evolving process of sense-making.
Acknowledgements
This piece was developed in close interaction with ChatGPT-4, used primarily as an assistant for structuring tables, refining grammar, compiling references, integrating citations upon request, and as an editor, improving the language. It did not originate core arguments, generate content independently, or challenge conceptual directions unless explicitly prompted to do so. At times, it served as a responsive surface—supporting, echoing, or helping me hold the thread of thought. Most importantly, ChatGPT-4, and later on GPT-5, served as the muse for writing this article.